OpenStudy (anonymous):

These convergence tests are slowly killing me. I need to find whether the following is absolutely convergent, conditionally convergent or divergent. And I can see it converges absolutely since it behaves like (2n^4)/(3n^6) but I can't figure out how to set up the comparison...and justify it

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty}(-1)^{n}\frac{ 2n^{4}-1 }{ 3n^{6}-n^{2}-1 }\]

OpenStudy (anonymous):

So the aim is to set something up where that < a convergent p-series, but the -n^2 -1 in the denom. is confusing me

OpenStudy (anonymous):

I realize that's what I'm supposed to do after comparing it to a p-series, it's just getting to it that's my problem D=

OpenStudy (rational):

\(n^6 \gt n^2+1\) for all \(n\gt 1\) \(\implies \dfrac{2n^4}{3n^6-(n^2+1)} \lt \dfrac{2n^4}{3n^6-(n^6)} \) for all \(n\gt 1\) agree ?

OpenStudy (anonymous):

Agreed and thank you!

OpenStudy (rational):

for \(n\gt 1\) we have \[\dfrac{2n^4-1}{3n^6-(n^2+1)} \lt \dfrac{2n^4}{3n^6-(n^2+1)} \lt \dfrac{2n^4}{3n^6-(n^6)} = \dfrac{1}{n^2}\] next you know what to do